Operation Summary for October 2, 2009 - October 5, 2009 |
The Integrated Luminosity for the period from 9/28/09 to 10/4/09 was 40.73 pb-1. Friday Oct. 2 The day shift began with Operations monitoring the Recycler stash (451.9E10), with the Antiproton Source stacking (123.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store. Operations put the Tevatron (TeV) into a dry squeeze at 8:14 AM followed by a wet squeeze. Operations put the TeV into shot setup at 10:25 AM. At 10:50 AM, the TeV quenched due to a Time Line Generator (TLG) problem. At 12:15 PM, the TeV quenched during a ramp due to a cryo system problem at TeV sector E3. At 3:17 PM, there was an indication of a TeV quench due to a power supply failure (C:AQ7), but operators investigated and found the low beta magnets still on. TeV personnel investigated further. Power supply personnel worked on the Linac Neutron Therapy Facility (NTF) 58˚ bend magnet power supply during the day shift. There were short downtimes during the day shift due to problems with RF stations in Linac, Booster, and MI. Operations reported that TeV support personnel completed their work on C:AQ7 at 5 PM. Operations completed dry and wet squeezes and orbit checks by 6:30 PM. Operations put the TeV into shot setup and established store 7237 at 10:15 PM with an average initial luminosity (AIL) of 251.71E30. At 11 PM, Operations reported that the Pbar 4-8 stochastic cooling system began tripping off. Operations called in Pbar personnel. Saturday Oct. 3 The midnight shift began with Operations monitoring store 7237 (135.74E30) and the stash (277.8E10), with Pbar stacking (26.4E10), and with NuMI and MiniBooNE taking beam. At 3:50 AM, Operations reported that a power supply (E:V118), the critical device for NuMI, had tripped off on an excessive ripple indication. Operations called in power supply personnel. Power supply personnel reported at 5:12 AM that they found a couple of bad capacitors for V118. Operations terminated store 7237 at 11 AM. Operations put the TeV into shot setup and established store 7238 at 1 PM with an AIL of 231.7E30. During the day shift EE support personnel continued work on the NTF bend magnet power supply. At 6:30 PM, a Pbar Accumulator quadrupole magnet (A6Q12) developed a LCW leak. Operations prepared Pbar for an access and called in water works personnel. Water works personnel completed their repair on A6Q12 magnet and power supply at 9:20 PM. Sunday Oct. 4 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. At 5:24 AM, Operations reported that the Booster West Anode Power Supply and a Linac quadrupole power supply (L:QPS104) tripped off. Everything reset. At 5:37 AM, a Tevatron Electron Lens (TEL) tripped off on a high vacuum indication. Operations terminated store 7238 at 8:18 AM. Operations put the TeV into shot setup and established store 7239 at 10:06 AM with an AIL of 220.73E30. Operations reported that the NuMI horn had tripped off numerous times throughout the day shift. It reset. Operations and CHL cryo system personnel reported a ramp permit problem for TeV sector C3. At 7:28 PM, CHL reported that they had lost helium compressor D. They brought the backup compressor online. The Cryo Coordinator reported at 11:16 PM that helium pressures had dropped, but they came back to normal when CHL personnel brought the backup compressor online. He also reported cryo personnel had fixed the ramp permit problems at TeV sector C3. Monday Oct. 5 The midnight shift began with Operations monitoring store 7239 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. Operations terminated store 7239 at 3:30 AM and put the TeV into shot setup. At 5:05 AM, the TeV quenched right after the antiprotons were loaded. The quench occurred at TeV sector F1 during flattop. Operations and TeV personnel investigated the cause. Power supply personnel resumed their work on the NTF bend magnet power supply at 7:56 AM. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow Recycler and MI studies, Booster and Linac coalescing work, and Switchyard and MI septa scans. |
|
More Information Comments and Suggestions |
| last modified 10/05/2009 email Fermilab |
| Security, Privacy, Legal | |